• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 06:29:16 on March18,2005

    Entry number 141805

    keyword=BB tdc spectra ok

    Now it looks like the BB tdc spectra are OK. Only some runs were problematic, for example 2759. ND tdc spectra are always OK. Suspecting the low tdc values in BB is due to near-to-dying VME, I did reset the ROC6 and did "startcoda" (probably I may not had to do that, just could have waited for next run to analyze). The run after this process (2761) shows good result.